satellite73 (satellite73):

anyone? \[f(x)=x^4+2x^3-4x^2-x+1\] let a,b,c,d be the roots of this ab, ac, ad, bc, bd, cd are roots of polynomial g(x) what is g(1) if g(0)=1

OpenStudy (anonymous):

this is what i tried so far, nothing came to mind http://openstudy.com/study?login#/updates/4f4d8bf0e4b019d0ebadd6e6

OpenStudy (anonymous):

hello zarkon thinking this would work out with vieta's formulas somehow but i cannot get it

OpenStudy (anonymous):

i tried a simpler example and nothing stood out. the way i understand it is we want the sum of the coefficients, but when i tried multipling out \[(1-ab)(1-ac)(1-ad)(1-bc)(1-bd)(1-cd)\] using worlfram, i thought i would see some nice pattern that would relate back to the elementary symmetric functions but it did not seem to work out that way

OpenStudy (zarkon):

I don't see a nice way to do this. I'll think about it though.

OpenStudy (anonymous):

I think ive figured out all the coefficients of g except for the x^3 and x^2. Its a real pain though, i dont know if I want to continue =/

OpenStudy (zarkon):

-9

OpenStudy (anonymous):

-9 c'mon now give me something to work with...

OpenStudy (anonymous):

hello joe!

OpenStudy (zarkon):

I'm sure there has to be a nicer way. I'm just doing a ground and pound....long tedious calculations using what we know about f and g. I get \[g(x)=x^6+4x^5-3x^4-13x^3-3x^2+4x+1\] I'm betting there is a slick way of doing this without capturing all the the coefficients of g. I just don't see it right now.

OpenStudy (anonymous):

so i am assuming you found all the zeros of f somehow and then multiplied to get this?

OpenStudy (zarkon):

Initially yes...but I have since determined the coefficients of g bases on the expansion of (x-a)(x-b)(x-c)(x-d) and abcd=1.( though I used my calculator to do all the expanding and factoring) The process is quite inelegant and I'm hoping that someone else has a more sophisticated solution.

OpenStudy (anonymous):

where is a -9 in \[f(x)=x^4+2x^3-4x^2-x+1\] minus the sum of the absolute value of the coefficients?

OpenStudy (zarkon):

I don't know.